OpenStudy (anonymous):

Can somebody tell me the solution for X. 2(x + 3)^2 = 5(x + 3) - 2

OpenStudy (amistre64):

quadratics are fun ... make a cleanup variable: a = x+3, so x = a-3 2a^2 = 5a - 2 ; when we know a we know x

OpenStudy (amistre64):

you might either complete the square, or use the quadratic formula

OpenStudy (amistre64):

a stands for x+3, its just a way to clean up the equation. solving for a is a simpler process than trying to square and distribute and gather like terms is all

OpenStudy (amistre64):

if we let a = x+3, then x = a-3 solving for a, gives us a way to find x

OpenStudy (anonymous):

-5/2 ?

OpenStudy (amistre64):

reworking this into a more proper format ... 2a^2 = 5a - 2 2a^2 -5a +2 = 0 either way you go you will need to either complete the square of run the quadratic formula, this is just bypassing alot of extra work

OpenStudy (amistre64):

5 +- sqrt(25-4(4)) a = ----------------- 4 5 +- sqrt(9) a = ----------- 4 a = 8/4 or 2/4 therefore; x+3 = 8/4 or 2/4

OpenStudy (anonymous):

I kinda understand it now. Thank you.

OpenStudy (amistre64):

youre welcome, and yeah: -5/2 i believe is one solution for it

OpenStudy (amistre64):

1/2 - 3 = -2 1/2 ... -5/2 2 - 3 = -1 is another solution